Let S1 be the part of the paraboloid z = 1 − x ^2 − y ^2 that lies above the plane z = 0. Let S2 be the part of the cone z = x ^2 + y ^2 + 2(sqrt till y^2) that lies inside of the cylinder x ^2 + y^ 2 = 1. Let S3 be the part of the cylinder x ^2 + y ^2 = 1 that lies between these surfaces. If S is the union of these three surfaces, find the flux over S of the vector field.

F = < cosz62 (yz), y + xz, 7x^ 2 + e^ tan(y)>
Answer #1

weve F = 200 S76292), 7+ x2, 7x + € 2. Jand to find flux over s by using div. thm which states that. SS F-05 = SSS divCF) av6 also weve x?ty? Let 20 = T10SO , y=rsino & dx dy = r drdol .. 3 ; 1+2 b 3-2 0 0 80%=8t2 161-82) Flux = 5PSs razardo Salon